### v2

# Autonomous Navigation of Quadrupeds Using Coverage Path Planning with Morphological Skeleton Map

Abstract
[…]
Index Terms:
Quadruped, Autonomous Navigation, Unstructured Environment, Coverage Path Planning, ROS 2
I Introduction
Due to advancements in technology and miniaturization, in the past decade surface (or ground) robots, such as wheeled and legged robots, have been increasingly adopted for diverse operations in harsh and unstructured environments. One of the key challenges in such environments is that the infrastructure to support diverse operations does not readily exist. These environments include, for example, disaster response [1, 2, 3], mining operations [4, 5], space exploration [6, 7, 8, 9], surveillance in remote locations [10, 11], or hazardous industries like nuclear power plant maintenance [12, 13].
In such complex environments, legged robots are more versatile and robust compared to wheeled robots than other surface robots such as wheeled rovers, and can adaptively navigate uneven, rugged, or soft terrain. Legged robots can cover relatively larger spatial areas by choosing safe footholds within their range of motion and rapidly responding to adjust their kinematic configuration [14] to achieve their objectives. The number of legs in a legged robot determines its movement efficiency and ability to maintain stability [15]. Compared to bipedal humanoids, quadruped robots demonstrate a greater load capacity and improved stability due to their broader base of support. On the other hand, quadrupeds possess simpler structures and control mechanisms than hexapodal and octopodal robots [16, 17]. For this reason, quadruped robots are ideal for tasks involving safe navigation of complex 3D environments for (sub-)surface exploration.
(a) Boston Dynamics Spot.
(b) Anybotics Anymal.
(c) Unitree Robotic Go2 Edu.
Figure 1: Commercially available quadrupedal robots from different companies.
There are several quadrupedal robots that are already commercially available on the market. We compare three notable examples, Boston Dynamics’ Spot, ANYbotics’ ANYmal, and Unitree Robotic’s Go2 Edu as seen in Figure 1, regarding attributes related to access of development, operation durability, and affordability, as seen in Table I. Both Spot and Anymal have garnered significant popularity and have made substantial contributions in research and engineering [18, 19] compared to that of Go2 Edu. However, their operational runtime is limited, and their cost is considerably higher. Go2 has three modes: Air, Pro, and Edu with a cost of 1600 USD, 2800 USD, and 12500 USD, respectively. However, only the Edu mode allows for software development, which is necessary for custom implementations, including other necessary features. Furthermore, Go2 Edu has a dedicated Robotics Operating System 2 (ROS 2) integration that allows rapid development and testing. With a factor of two to three in running time, experiments can be done over a long session and the robot can provide sufficient battery capacity to power additional sensors. This paper focuses on the development of a software architecture for path planning and navigation specific to Go2 Edu.
The goal of this research is to enable quadruped robots to map the terrain using coverage path planning. For quadrupeds to do this, they require sensors, where the most common are cameras [20, 21, 22] and LiDAR [23, 24]. To be able to scan with the sensors, the robot needs to move to various POIs, which requires coverage path planning and navigation methods. These will be examined in the related works below.
Table I: Comparison of commercially available quadruped robots.
Feature Spot ANYmal D Go2 Edu
Manufacturer Boston Dynamics ANYbotics Unitree Robotics
Dimensions (L/W/H, mm) 1100 x 500 x 191 930 x 530 x 890 700 x 310 x 400
Maximum walking speed (m/s) 1.6 1.3 3.7
Average running time (min) 90 90-120 120-240
Integrated LiDAR No Yes Yes
Integrated optical camera Yes Yes Yes
Integrated depth camera Yes Yes No
Connectivity Wi-Fi, Ethernet Wi-Fi, 4G Wi-Fi, Bluetooth, 4G, Ethernet
Custom Software Development Supported (SDK, APIs) Supported (ROS Integration, APIs) Supported (SDK, ROS Integration)
Estimated cost (1000 USD) 74.5 150 12.5
I-A Related Works
Several studies have recently explored the development of software infrastructure for the Unitree Robotics Go2 Edu robot. In [25], a reinforcement learning method was developed to enable the Go2 Edu robot to navigate narrow pipes using visual inputs from a depth camera. The navigation process is relatively straightforward due to the grid-like structure of the pipes, where each pipe is aligned in a straight path with occasional protrusions serving as obstacles. The work in [26] developed a high-level path planning using a large language model, namely OpenAI’s ChatGPT-4o, which allows the interpretation of human verbal commands and translates them into a list of executable instructions. The system integrates a depth camera alongside a segmentation model to perceive the environment effectively. In addition, research in [27] developed a motion controller for the Go2 Edu robot to traverse complex and unstructured environments using proprioceptive sensing and collision estimation only.
Autonomous navigation using quadrupedal robots is crucial for exploring complex environments; however, limited work has been done on 2D coverage path planning. Research in [28] achieved coverage path planning for loco-manipulation through an integrated end-to-end pipeline combining perception, optimization, and whole-body motion planning with RGB-D camera inputs. Another work in [23] presented a 2D coverage path planner for investigating unknown and unstructured environments while accounting for time-bounded and dynamics constraints, and traversability risk. The advantage of these approaches is that they guarantee timely execution when the mission is time-bound, and they find a good optimal tradeoff between maximum area coverage and path traveled. The drawback of that approach is that it is focused on time constraints, therefore, it does not work when the task has a variable execution time and requires time to be defined in advance. In contrast, our approach enables planning for variable times.
Some navigation methods already enable planning for variable times. In the recent work [24], a novel autonomous exploration method was developed using a topological skeleton of the environment’s geometry via LiDAR and a finite state machine (FSM) that enables exploratory strategy. This was demonstrated on a quadrupedal robot in an unstructured environment. The advantage of these approaches is very adaptable in an unknown environment due to the use of a state machine that continuously checks for undiscovered areas within the map. The main difference between this work and our approach is that their method for obtaining topological skeletons uses wave propagation and Voronoi diagrams, while our method uses a morphological technique by treating the environment as an image. The impact of this is that the morphological technique simplifies the mapping and is thus computationally more efficient.
Furthermore, all the above methods on coverage path planning have specific map generation algorithms that are an integral part of the whole approach. Therefore, they are less modular, and it is not easy to replace them with developing state-of-the-art mapping algorithms. In contrast, our approach can work with different types of mapping algorithms, thus making it more modular.
I-B Contribution
In this work, we address the gap in the state-of-the-art with the following key contributions.
  1. 1.
    We develop a control framework which is based on a finite state machine that can switch between different operation modes that enable autonomous navigation and environmental inspection.
  2. 2.
    Using a prior 2D navigation map of the surroundings, the framework rapidly generates an efficient path based on the morphological skeleton of the map, which ensures coverage from a small to a large area.
  3. 3.
    We validate the developed system with on-field experiments on navigational tasks using the Unitree Robotics Go2 Edu robot in an indoor environment.
  4. 4.
    We designed an extended interface that enables a Unitree Robotics Go2 Edu to work together with the control framework using ROS 2.
  5. 5.
    The whole application is open-source, and it is available at https://github.com/asil-lab/go2-autonomous-navigation
II Preliminaries
II-A Hardware Specifications
The Unitree Robotics Go2 Edu features three degrees of freedom (DOFs) per leg, consisting of hip, thigh, and calf hinge joints (from base to foot). It is equipped with an inertial measurement unit (IMU), a HD wide-angle camera, and foot-end force sensors. The robot offers a battery life of 2 to 4 hours and supports fast charging  [^1] Go2 SDK Development Guide - About Go2 https://support.unitree.com/home/en/developer/about_Go2.
For navigation and perception, the Go2 Edu is fitted with the Unitree L1- a 4D LiDAR (3D position + 1D greyscale) based on laser time-of-flight (TOF), mounted on its mouth. This LiDAR provides a 360° × 90° field of view (FOV), a measurement accuracy of ±2.0 cm, and a scanning distance of up to 30 m with 90% reflectivity. It integrates an IMU with a 3-axis accelerometer and 3-axis gyroscope, has a proximal blind spot of 0.05 m, and features a sampling frequency of 43,200 points per second. Additionally, it operates with a circumferential scanning frequency of 11 Hz and a vertical scanning frequency of 180 Hz 2224D Lidar L1 Application Scenarios 4D Lidar L1 Efficacy — Unitree Robotic https://www.unitree.com/LiDAR. The integrated LiDAR sensor is also utilized for collision detection which facilitates the differentiation between free and occupied spaces in the environment with respect to its height. Additionally, it enables the generation of a 2D map for navigation during run-time, and self-localization ability with respect to environmental features and the current state of the map.
Furthermore, the robot includes an expansion dock, which houses an NVIDIA Jetson Orin with a computing power of 40–100 TOPS. It also comes with a manual two-handed joystick controller for user operation. Additionally, in order to monitor the robot remotely, a TP-Link TL-WR802N Nano WLAN Router is added to set a wireless connection between the robot and the operator’s computer.
Lastly, we also integrated external sensors on the robot to measure ambiance such as temperature, humidity, and light intensity. The integration of these sensors is discussed in Appendix -A, but the usage of these sensors is outside the scope of this work and is hence not detailed in this paper.
II-B Software Specifications
The Unitree Robotics Go2 Edu robot has a dedicated Software Development Kit (SDK) which allows custom implementations to be programmed. This SDK uses a Data Distribution Service (DDS) as the networking middleware, which enables reliable and real-time data exchange between the program and the robot  [^3] Go2 SDK Development Guide - SDK Concepts https://support.unitree.com/home/en/developer/SDK_Concepts.
Using DDS, a Robotics Operating System (ROS) application [29] is also implemented to facilitate seamless communication between distributed robotic components, ensuring real-time data exchange, scalability, and interoperability across diverse hardware and software platforms.
In particular, we use ROS 2, given the additional benefits as compared to ROS 1 such as decentralization, simplicity, user-friendliness to name a few. We use the ROS 2 version Foxy on the Linux Ubuntu OS version 20.04 to develop our proposed framework, since those specifications are well established in the Unitree Robotics Go2 Edu robot.
II-C SLAM
Figure 2: The pipeline diagram of filtering the point cloud for SLAM and 3D scanning per measurement update.
In order for the robot to determine its location within the environment while simultaneously creating a map, we use Simultaneous Localization and Mapping (SLAM). SLAM allows the robot to dynamically create a map based on the history of information, and localize the robot as a function of both the current measurement and the map, simultaneously [30]. In our case, a 2D map is sufficient because the robot can only navigate on the ground surface in $x$- and $y$-direction (excluding orientation in the $z$-direction). It is worth mentioning that Go2 Edu already comes with its own SLAM implementation. However, at the time of the development, it was not readily available due to lack of documentation. Moreover, we would also like to keep modularity in mind such that this implementation is applicable to other quadrupeds besides the software modules bridging the proposed method and the robot. Therefore we can use Macenski’s SLAM Toolbox to create a 2D map using a sequence of 2D laser scans as input [31]. This 2D laser scan is the measured distance reflected against an obstacle with respect to the robot’s laser scanner at some angle in some given time. Using a radial laser scanner, it can measure the layout of surroundings around the robot simultaneously in one measurement timestep.
Despite Macenski’s impressive SLAM implementation, the integrated sensor that measures the geometry of the surroundings is a LiDAR that outputs the measurements as a point cloud in 3D. To make it interpretable for SLAM Toolbox, we process these LiDAR measurements to identify the obstacles which the robot cannot navigate through. To achieve this task, we develop a data processing pipeline using the Point Cloud Library (PCL) [32] to transform the input 3D point cloud /point _cloud/raw in order to obtain the desired output 2D laser scans /scan. An overview of this pipeline is shown in Figure 2. Note that this pipeline also outputs another point cloud /point_cloud/sampled, which will later be used for environment scanning described in Appendix -C.
II-D Navigation
We use the Nav2 framework [33] to ensure the robot is able to plan and navigate towards a desired pose (position and orientation) in the environment, which is also completely compatible with the SLAM Toolbox discussed previously. Using the 2D navigation map provided by SLAM, the Nav2 framework can plan and navigate a path as a function of the destination’s pose, the robot’s current pose, and the robot’s kinematics constraints with respect to its surroundings. A 2D pose is defined as $\boldsymbol{x}:=\begin{bmatrix}x&y&\psi\end{bmatrix}^{T}$, where $x$ and $y$ define the longitudinal and lateral displacements, and $\psi$ as the orientation of the robot about the $z$-axis (vertical displacement).
The Nav2 framework includes an array of tools such as planners, recoveries and controllers. It is outside the scope of this paper to experiment with different tools. Therefore, we mainly used the default configuration with minor adjustments that corresponds to the robot’s kinematics. Due to the fact that the robot’s movement can be controlled using 2D velocities $\dot{x},\dot{y},\dot{\theta}$ as inputs, we can consider the robot to behave similarly to that of a differential wheeled robot. With this in mind, we can simply use Nav2’s default planner NavFnPlanner.
Lastly, since the Nav2 framework requires a desired pose as an input, it serves as a local planner and navigation. Therefore, in order to achieve autonomy in the robot, we require a higher-level navigation approach capable of identifying and selecting regions of interest (ROIs) to navigate within the environment.
III Methods
In this work, we propose a novel framework for the Unitree Robotics Go2 Edu robot for navigation in unstructured and unpredictable environments. The diagram in Figure 3 shows an overview with key modules. The key modules that are responsible for the autonomous navigation are highlighted in red. To create a graph of ROIs, the map reader looks at the SLAM-generated map and transforms it into a topological skeleton based on its geometry (Sec. III-A). The path planning is informed about the graph which creates an efficient path of ROIs, referred to as waypoints, that the robot must take during navigation (Sec. III-B). To control the high-level operation, a state machine (Sec. III-C) is used to switch between actions, e.g., checks when/if each waypoint is complete, if a fallback strategy is needed, if human operator input is detected, etc. The communication between the modules is done via ROS 2 using an extended interface. More information on the extended interface is found in Appendix -B.
Figure 3: The system overview of the proposed autonomous navigation with scanning capabilities in an unstructured environment. Green: hardware modules, blue: SLAM and Nav2 modules, red: autonomous navigation modules, purple: supplementary modules, and yellow: human agent.
III-A Map Reader
To effectively cover the environment, the robot must follow a path that ensures it traverses every corner of the space. This path should align with the trajectory of the occupied areas, e.g. corridors. This approach is generally effective under the assumption that the environment is confined within a bounded, occupied space, such as an indoor setting. In order to create the path, we process the 2D navigation map as an 8-bit array using the algorithm as seen in Algorithm 1. $\mathbf{M}$, $H$, $W$ denote the map, and the map’s height and width, respectively. The map is divided into cells using the map resolution $R$ in m/pixel. Each of these cells is an element in $\mathbf{M}$, $m_{ij}$, which only contains one of the three types of space: occupied, free, and $unknown$, which are represented by the integers 0, 255, and 128, respectively.
Algorithm 1 Process the 8-bit 2D navigation map $\mathbf{M}$ into an unordered set of waypoints $\mathcal{V}$ where each element is a waypoint $\boldsymbol{v}\in\mathbb{R}^{2}$.
1:Input: Navigation map i.e., $\mathbf{M}\in\mathbb{N}^{H\times W}$
2:Input: Hyperparameters $\sigma,\kappa,\mathbf{K},\boldsymbol{o}\in\mathbb{R}^{2},R$
3:$\mathcal{V}\leftarrow\emptyset$
4:for $i\in\{1,\ldots,H\}$ do
5:     for $j\in\{1,\ldots,W\}$ do
6:         if $[\mathbf{M}]_{ij}<255$ then
7:              $[\mathbf{M}]_{ij}=0$
8:         end if
9:     end for
10:end for
11:$\mathbf{M}\leftarrow$ GaussianFilter($\mathbf{M}^{\prime}$, $\sigma$)
12:for $i\in\{1,\ldots,H\}$ do
13:     for $j\in\{1,\ldots,W\}$ do
14:         if $[\mathbf{M}]_{ij}>\kappa$ then
15:              $[\mathbf{M}]_{ij}=1$
16:         end if
17:     end for
18:end for
19:$\mathbf{C}\leftarrow\arg\max_{|\boldsymbol{c}_{i}|}$ FindContours($\mathbf{M}$)
20:$\mathbf{M}\leftarrow$ FillAreaByContour($\mathbf{C}$)
21:$\mathbf{M}\leftarrow$ Erode($\mathbf{M}\mid\mathbf{K}$)
22:$\mathbf{M_{\mathbf{S}}}\leftarrow$ Skeletonize($\mathbf{M}$)
23:for $i\in\{1,\ldots,H\}$ do
24:     for $j\in\{1,\ldots,V\}$ do
25:         if $[\mathbf{M}]_{ij}=255$ then
26:              $\mathcal{V}\leftarrow\mathcal{V}\cup\{R\begin{bmatrix}i&j\end{bmatrix}^{T}+\boldsymbol{o}\}$
27:         end if
28:     end for
29:end for
30:return $\mathcal{V}$
First, consider an indoor environment. In such case, most unknown cells inside the map are located outside the boundaries of the occupied area e.g. beyond walls. This is because SLAM initially represents the environment as a grid of unknown cells before any exploration occurs. Furthermore, unknown cells often persist between occupied and free regions as a result of occlusions encountered during raycasting-based measurements, such as those performed using LiDAR, as seen in Figure 6(a). Consequently, this unknown space cannot be reached by the robot and we can simply treat the unknown regions as part of the boundaries–the occupied space. This can be seen in Figure 6(b).
To establish a smooth connection between occupied cells affected by noise in the map, a Gaussian filter [34] is applied using a standard deviation parameter $\sigma\in\mathbb{N}$, as seen in Figure 6(c). Subsequently, in order to restore the binary representation of occupied and the free cells, the filtered result is binarized using a threshold parameter $\kappa\in[0,255]$. A cell with a value greater than $\kappa$ is classified as a free cell; otherwise, it is designated as an occupied cell.
Due to residual noise and the presence of transparent obstacles, such as windows, this means that there may be free cells that is unreachable to the robot. To mitigate this issue, we assume that the navigable space for the robot corresponds the largest 2D contour. First, the contours within the map are found using the marching cubes algorithm [35]. The map is then reconstructed by filling the largest 2D contour with a value of 255, as seen in Figure 6(d). It is worth noting that by only filling the largest contour, it does not take obstacles inside the contour into account.
To ensure that the robot remains a safe distance from the occupied space during navigation, the free space in the map is shrunk using a morphological filter called erosion [36], as seen in Figure 6(e). Erosion uses a structuring element $\mathbf{K}$, also referred to as a kernel, that determines the width to be removed. For the sake of simplicity, we take a square matrix of ones as our structuring element: $\mathbf{K}=\boldsymbol{1}_{k}\boldsymbol{1}_{k}^{T}$, where $k$ determines the length of the vector and $k=1,2,3,\ldots$. The larger the $k$, the larger the safe distance. In addition, the safe distance reduces when $R$ reduces. Therefore, $k\propto R$. This also removes narrow passages that prove impassable for the robot.
The remaining free space is reduced to a thin one-pixel-wide representation containing $O$ pixels, which corresponds to the topological skeleton of the map’s geometry. This can be seen in Figure 6(f). In a practical scenario, the time complexity of 2D skeletonization is mainly proportional to the total number of pixels in an image $P$ as it iterates until the object becomes one-pixel wide. Therefore, it can be considered as $\mathcal{O}(P)$ [37].
The 2D skeleton map $\mathbf{M_{\mathbf{S}}}$ is then flattened into an unordered set that describe the $x$- and $y$-position of every 2D point by transforming the pixel coordinates into real-world coordinates using the origin $\boldsymbol{o}$ and resolution $R$ of the 2D navigation map. Only those pixels for which $m_{\mathbf{S},ij}=255$ are considered in this transformation. Consider that the skeleton contains $O$ pixels, it follows that there are $O$ waypoints in $\mathcal{V}$. We refer to every such 2D point defined by the skeleton as waypoints, denoted as $\boldsymbol{v}_{i}\in\mathbb{R}^{2}$ where $0\leq i\leq O$. The waypoints serve as points of interests for the robot to visit.
III-B Path Planning
In order to scan the whole environment, the robot should perform the scanning procedure for each waypoint. Therefore, the objective is to determine a path that includes every waypoint the robot must visit while optimizing for time efficiency. Ultimately, this can be considered as a Travelling Salesman Problem (TSP) which we approach by formulating a fast and efficient path planning algorithm. It takes a connected acyclic graph $\mathcal{G}$ as input, and outputs a path $\mathcal{P}$, which is the ordered set of waypoints. In other words, $\mathcal{P}$ is a sequence of vertices $\boldsymbol{v}_{i}$ traversed by the robot.
In order to construct the graph $\mathcal{G}$, we treat the unordered set of waypoints $\mathcal{V}$, obtained from the map reader, as the set of vertices. Each vertex in $\mathcal{V}$ uniquely corresponds to the coordinate vector of a waypoint, and can, for the sake of simplicity, be denoted as $\boldsymbol{v}_{i}\in\mathbb{R}^{2},\;\forall i=1\ldots,O$. Each vertex is then connected to other vertices if they are neighbors around the map’s resolution $R$. This should resemble to the skeleton representation of the map, where the connections define the edges $\mathcal{E}$ of the graph $\mathcal{G}$.
1:Input: Define graph $\mathcal{G}(\mathcal{V},\mathcal{E})$
2:Hyperparameters: $R$, $D$, $\boldsymbol{x}_{0}$
3:$\mathcal{V}_{\text{leaf,total}}\leftarrow\emptyset$
4:for $\boldsymbol{v}_{i}\in\mathcal{V}$ do
5:     if degree$(\boldsymbol{v}_{i})=1$ then
6:         $\mathcal{V}_{\text{leaf,total}}\leftarrow\mathcal{V}_{\text{leaf,total}}\cup\{\boldsymbol{v}_{i}\}$
7:     end if
8:end for
9:$\mathcal{V}_{\text{leaf,visited}}\leftarrow\emptyset$
10:$\mathcal{P}\leftarrow\emptyset$
11:$\boldsymbol{v}_{\text{leaf,start}}\leftarrow\text{FindNearestLeaf}(\boldsymbol{x}_{0},\mathcal{V}_{\text{leaf,total}})$
12:while $|\mathcal{V}_{\text{leaf,visited}}|<|\mathcal{V}_{\text{leaf,total}}|$ do
13:     $\mathcal{V}_{\text{leaf,visited}}\leftarrow\mathcal{V}_{\text{leaf,visited}}\cup\{\boldsymbol{v}_{\text{leaf,start}}\}$
14:     $\boldsymbol{v}_{\text{leaf,target}}\leftarrow\text{FindNearestLeaf}(\boldsymbol{v}_{\text{leaf,start}},\mathcal{V}_{\text{leaf,total}}\setminus\mathcal{V}_{\text{leaf,visited}})$
15:     $\mathcal{P}^{\prime}\leftarrow\text{FindShortestPath}(\boldsymbol{v}_{\text{leaf,start}},\boldsymbol{v}_{\text{leaf,target}}\mid\mathcal{G})$
16:     for $\boldsymbol{v}^{\prime}_{i}\in\mathcal{P}^{\prime}$ do
17:         if $\boldsymbol{v}^{\prime}_{i}\not\in\mathcal{P}$ then
18:              $\mathcal{P}\leftarrow\mathcal{P}\cup\{\boldsymbol{v}^{\prime}_{i}\}$
19:         end if
20:     end for
21:     $\boldsymbol{v}_{\text{leaf,start}}\leftarrow\boldsymbol{v}_{\text{leaf,target}}$
22:end while
23:$\mathcal{P}^{*}\leftarrow\emptyset$
24:for $k\leftarrow 0$ to $\left\lfloor{\frac{|\mathcal{P}|-1}{D/R}}\right\rfloor$ do
25:     $\mathcal{P}^{*}\leftarrow\mathcal{P}^{*}\cup[\mathcal{P}]_{\frac{D}{R}k}$
26:end for
27:return $\mathcal{P}^{*}$
Algorithm 2 A fast and efficient approach to planning a path $\mathcal{P}^{*}$, with $L$ number of vertices, using the unordered set of waypoints obtained from map reader $\mathcal{V}$, map resolution $R$, waypoint resolution $D$, and the robot’s starting 2D position $\boldsymbol{x}_{0}$.
Algorithm 2 outlines the procedure for obtaining an efficient path $\mathcal{P}^{*}$ for coverage exploration of the whole environment. First, there are dead-ends found in the skeleton representation of the map. For the robot to cover the whole environment, we can utilize these dead-ends. Each dead-end is identified as a leaf vertex $\boldsymbol{v}_{\text{leaf}}\in\mathcal{V}_{\text{leaf}}$ and is typically defined with degree 1. Assuming the graph $\mathcal{G}$ is connected and acyclic, complete coverage of the map can be achieved by ensuring that the robot visits every leaf vertex in $\mathcal{V}_{\text{leaf}}$ in sequence. Since all other vertices in the graph lie on the paths connecting the leaves. traversing to each leaf inherently requires passing through the intermediate vertices. As a consequence, visiting all leaf vertices implies that the entire graph has been traversed.
After identifying all leaf vertices, the leaf vertex closest to the robot’s current 2D position $\boldsymbol{x}_{0}$ is selected as the initial source leaf vertex $\boldsymbol{v}_{\text{leaf,start}}$ using Algorithm 3. For every source leaf vertex $\boldsymbol{v}_{\text{leaf,start}}$, we find the next nearest unvisited leaf vertex as the target $\boldsymbol{v}_{\text{leaf,target}}$ using FindNearestLeaf, as defined in Algorithm 3. We find the shortest path $\mathcal{P}^{\prime}$ between the source $\boldsymbol{v}_{\text{leaf,source}}$ and the target $\boldsymbol{v}_{\text{leaf,target}}$ given the graph $\mathcal{G}$. To achieve this, we use the method FindShortestPath which essentially utilizes Dijkstra’s algorithm. This algorithm has a time complexity of $\mathcal{O}(|\mathcal{E}|+|\mathcal{V}|\log|\mathcal{V}|)$, where $|\mathcal{V}|$ and $|\mathcal{E}|$ denote the total number of vertices and the total number of edges in the graph, respectively [38].
Once $\mathcal{P}^{\prime}$ is found, each vertex $\boldsymbol{v}^{\prime}_{i}\in\mathcal{P}^{\prime}$ is appended into $\mathcal{P}$. Unless some vertex $\boldsymbol{v}^{\prime}_{i}$ reappears as a path of a different path $\mathcal{P}^{\prime}$, revisiting and scanning at this position is unnecessary and should be avoided to conserve time and computational resources. Once iterated over all $\boldsymbol{v}^{\prime}_{i}$ in $\mathcal{P}^{\prime}$, $\boldsymbol{v}_{\text{leaf,start}}$ is inserted into a set of visited leaf vertices $\mathcal{V}_{\text{leaf,visited}}$ and $\boldsymbol{v}_{\text{leaf,target}}$ becomes the next source $\boldsymbol{v}_{\text{leaf,start}}$. Not only is the shortest path useful for appending $\boldsymbol{v}_{i}$ into $\mathcal{P}$ but also for ensuring a feasible and unobstructed trajectory through occupied spaces.
1:Function FindNearestLeaf
2:Input: $\boldsymbol{x}_{0}\in\mathbb{R}^{2}$, $\mathcal{V}^{\prime}_{\text{leaf}}$
3:$\boldsymbol{v}_{\text{leaf,nearest}}\leftarrow\boldsymbol{0}$
4:$d\leftarrow\infty$
5:for $\boldsymbol{v}_{\text{leaf,i}}\in\mathcal{V}^{\prime}_{\text{leaf}}$ do
6:     if $|\boldsymbol{v}_{\text{leaf,i}}-\boldsymbol{x}_{0}|<d$ then
7:         $\boldsymbol{v}_{\text{leaf,nearest}}\leftarrow\boldsymbol{v}_{\text{leaf,i}}$
8:         $d\leftarrow|\boldsymbol{v}_{\text{leaf,i}}-\boldsymbol{x}_{0}|$
9:     end if
10:end for
11:return $\boldsymbol{v}_{\text{leaf,nearest}}$
12:End Function
Algorithm 3 A method to find the nearest leaf vertex $\boldsymbol{v}_{\text{leaf,target}}$ from a set of selected leaf nodes $\mathcal{V}^{\prime}_{\text{leaf}}$ to a given source $\boldsymbol{x}_{0}$ given the Euclidean distances.
This is iterated until all leaf vertices have been added and $\mathcal{P}$ is complete. Nevertheless, because the original distance between every two waypoints is approximately equal to the map’s resolution, this leads to significant time spent on scanning with minimal displacement. Considering that the robot has a large scanning range capability of 30 m, which allows it to scan distances greater than the map resolution, we can increase the distance between every two waypoints to some arbitrary distance $D$ by reducing $P$ by a factor of $D/R$ where $D\geq R$. This is done by selecting every $D/R$-th element in $\mathcal{P}$.
III-C State Machine
(a) The activity diagram describing the autonomous coverage path planning given a prior 2D navigation map.
(b) A FSM diagram based on the activity diagram, where the decision arguments $i$ is the number of remaining waypoints to visit, $\mathbf{x}$ is the robot’s current 2D pose, $\mathbf{x}_{\text{des}}$ is the destination’s 2D pose, and $\boldsymbol{\delta}$ is the predefined tolerance.
Figure 4: Comparison of (a) the activity diagram for autonomous coverage path planning and (b) its corresponding FSM diagram.
To achieve autonomy that enables the robot to visit the waypoints and perform a scanning procedure in succession, an activity diagram is formulated as seen in Figure 4(a). We can group two or more activities as a singular state if the transitions are consecutive without interruptions, e.g., a decision node. This state will perform all activities in their respective order. On the other hand, each decision node becomes a state that checks if its control variable has reached a specified threshold. Each state has a conditionless trigger that automatically transitions the current state to the next specified state at the end of its action.
With a finite state machine (FSM), we can achieve the desired autonomous navigation based on the triggers by handling the sequence of actions. The implemented state machine diagram can be seen in Figure 4(b). Each state is explained as follows:
  • In State Load Map, an existing 2D navigation map is loaded onto SLAM and allows the robot to be localized with respect to the map’s metadata (meter-per-pixel resolution, width and height both in pixels, and origin as $(x,y)$ in meters), and the robot’s concurrent surroundings through 2D laser scans. After this map is loaded, it is then analyzed to generate a list of waypoints, each in $(x,y)$ coordinates, for the robot to visit using the map reader. An optimal route is then planned to visit all of these waypoints, which reorders the original list using the path planner. Finally, it automatically transitions to the next state, Check Waypoints.
  • State Check Waypoints allows the system to iterate over the list of waypoints. If there are waypoints remaining, it removes the waypoint in the first entry of the list and stores it as the current destination, and transitions to State Check Destination. If there is no waypoint left, it transitions to State Home.
  • In State Check Destination, the state machine determines if the robot is already at the current destination with some acceptable offset. This is done by determining if the Euclidean distance between the robot’s 2D pose $\boldsymbol{x}$ and a destination’s pose $\boldsymbol{x}_{\text{des}}$, is less than or equal to a set tolerance $\boldsymbol{\delta}$, where $\boldsymbol{x},\boldsymbol{x}_{\text{des}},\boldsymbol{\delta}\in\mathbb{R}^{3}$ since the poses are defined in the $x$- and $y$-axes, and the yaw orientation $\psi$ with respect to /map frame. If the condition is true, it transitions to State Scan. Otherwise, it transitions to State Move. It is worth noting that the path planner does not account for the orientation $\psi$. However, this orientation is likely a necessary requirement of the scanning procedure, ensuring that the robot is aligned with the desired orientation, as described in Supplementary Material C.
  • In State Move, it actuates the robot to navigate towards the desired destination. Navigation is considered successful when the Euclidean distance between the robot’s current pose $\boldsymbol{x}$ and the target destination $\boldsymbol{x}_{\text{des}}$ is less than or equal to a predefined threshold $\delta$. Alternatively, if the navigation process exceeds a specified timeout duration $T_{\text{timeout}}$, it is also terminated. In either case, the system transitions back to the Check Destination state. However, if the system detects an interruption by a human operator via the joystick controller mid-operation, the system promptly cancels the ongoing action and immediately transitions to State Manual Control.
  • In State Scan, the robot performs the procedure to scan the local environment, which is detailed in Supplementary Material C. Once it is completed, it transitions to State Check Waypoints.
  • State Manual Control allows the human operator to take over the robot’s navigation and move it towards the destination. This should happen in a case when the robot is traversing difficult terrains, or when the navigation framework is stuck at finding the right solution. This transitions to State Scan once the operator presses a button on the controller.
  • In State Home, the robot travels back to its starting position. Once the robot arrives at the starting position, it lies down on the ground and waits for the new commands.
To sum up, the proposed framework for navigation consists of three key modules: the map reader to extract POIs using the 2D navigation map as waypoints, the path planner to order the waypoints as an efficient route given the robot’s current position, and the state machine to enable the robot to navigate towards every waypoints and scan, consecutively.
IV Experiments and Results
To demonstrate the proposed framework and evaluate how well it performs, we conducted the experiment in a leveled in-door obstacle-free non-convex environment, which emulated an inaccessible environment. For evaluation, we used the following two metrics:
  1. 1.
    Time efficiency: the time required to process the map to create a path, and to reach each waypoint consecutively.
  2. 2.
    Reachability: the number of waypoints that the robot is able to reach over the total number of waypoints planned in %.
The Go2 Edu robot was initially manually controlled via a wireless controller to generate a 2D navigation map using the SLAM module. The experiment was then conducted and the whole process was repeated over five trials. The SLAM module’s map resolution was set to $0.10$ meters. The point cloud buffer time was set to 0.25 seconds. The laser scan’s minimal and maximal range was set to 0.50 meters and 30.0 meters, respectively. The smoothing standard deviation was set to 3. The crispification threshold was set to 128. The erosion kernel $K$ was set to 10 x 10 pixels. The waypoint-to-waypoint distance $D$ in path planning was set at 1.00 m. The robot’s maximum x-, y- and, yaw velocities were set to 1.00 m/s, 0.50 m/s, and 0.80 rad/s, respectively. The 2D navigation position and orientation tolerance were set to 0.05 m and 0.08 rad, respectively, with a navigation timeout of 10.0 s to replan. The scanning procedure module was disabled to demonstrate the navigational tasks.
IV-A SLAM, Map Reader and Path Planning
Figure 5: derived by normalizing over the five 2D navigation maps obtained from five trials, where the inner corner of the left wing of the room is taken as the origin which is denoted as a red point. All five maps are translated and rotated around the red point accordingly.
Before presenting the main results, we first validate the accuracy of the 2D navigation maps generated by the SLAM module. The average map is derived by normalizing over the five 2D navigation maps obtained from five trials, as illustrated in Figure 5. The width and height of these maps range between $[196,225]$ and $[185,231]$ pixels, respectively. We take the inner corner of the left-wing (upper left triangle of the map) of the room as the origin to align the maps in position and orientation accordingly. Analysis of the averaged map reveals that the right-wing exhibits a wider range of shades of gray about the occupied space, suggesting that this part of the environment is slightly tilted relative to its actual orientation. Nonetheless, this does not severely affect the map reader, path planning, and navigation as the robot mainly localizes itself in its immediate surroundings and the topological skeleton of the map can still be found. For instance, as seen in Figure 6, the map reader can still create waypoints mainly using the map’s geometry where it also enables the robot to visit the corners by producing branches from the main path to these corners.
(a) The original map.
(b) The adjusted map.
(c) The fuzzied map.
(d) The contour map.
(e) The eroded map.
(f) The skeleton map.
Figure 6: A demonstration of the map reader’s pipeline using the 2D navigation map of Trial 5 as input.
To evaluate the time efficiency of the map reader, we record the duration required to process each map across five trials, considering a total of $N=100$ instances. We vary the weights in terms of the dimension per map. To facilitate a clearer representation in the plot, we use the product of the dimensions, $H\times W$ in pixels. As can be seen in Figure 7(a), we can see that the average time taken for the map reader ranges from 2.34 ms to 2.70 ms over the five trials with an average standard deviation of about 0.30 ms. We can also see that the mean time increases by 22.0 ns for every additional pixel in the map. For instance, with a map with size $H\times W=10^{6}$ pixels, it has an expected mean time $T_{\text{read}}\approx 23.6$ ms.
To remain within a maximum time taken of 1.0 s, the map should not be larger than the size of $H\times W=45.4\times 10^{6}$ pixels. Assuming a square map with a map resolution of 0.10 m, this maximum map has a length of 673.4 m. Hence, our proposed map reader is fast for relatively small to large map sizes (ranging up to $45.4\times 10^{6}$ pixels) within a lead time of $T_{\text{read}}=1.0$ s, especially when the map reader is only run once to produce the waypoints given the map’s geometry.
[…]
To stay within a lead time of 1.0 s, the number of waypoints should not be larger than $1.219\times 10^{5}$. The bottleneck appears if the map resolution $R$ becomes small, as it increases the number of pixels per meter in a map. This can substantially increase the number of waypoints by a factor of $R/R^{\prime}$, where $R^{\prime}$ is the new map resolution and $R^{\prime}<R$. Nonetheless, the path planner can be considered fast for a small to large number of waypoints, given a reasonable map resolution $R$. It is also worth noting that it also runs only once to produce the necessary path.
[…]
Figure 7: Comparison of mean time taken for (a) map reader and (b) path planner over five trials.
To verify the path planner’s performance, we first evaluate the logic whether it works as intended. Using one trial as an example, we can see that the path planner has created the optimal path with respect to time based on the waypoints given the robot’s current 2D position, as seen in Figure 8(a). Furthermore, according to Figure 8(b), we can see that the mean waypoint-waypoint distance is about $1.00$ m given the previously set waypoint resolution $D$. The variance can be explained due to the fact that robot skips already visited waypoints, since it is unnecessary for the robot to scan about the same position twice. This results in the robot having to travel distances of $1.00$ m or more for some intervals. In contrast to traveling longer distance, the robot can also visit the next waypoint in a shorter distance. This is due to the fact that the path splicing happens at the end of the algorithm, whereby resulting in a probability that two waypoints are less than the set waypoint resolution. Nevertheless, this does not heavily affect navigation, and the average waypoint distance is almost identical to the set waypoint resolution $D$.
(a) An efficient path to explore the whole environment from the robot’s starting position (green) to the farthest waypoint (blue) while visiting all other waypoints (black). Each arrow (red) denotes the direction from source to target.
(b) Mean distance between every two connected waypoints generated by the path planning per trial, as defined by the waypoint resolution $D$.
Figure 8: Demonstrations of path planner’s performance.
IV-B Navigational Performance
Table II: Reachability and time taken for each trial including key observations.
Trial Reachability (%) Total Time (s) Median Time per Waypoint (s) Observations
1 100.0 443.2 5.40 Human assistance required at Waypoint 17.
2 100.0 397.0 5.30 Human assistance required at Waypoint 36.
3 82.61 287.2 5.40 Significant map drift occurred after 3 minutes, such that the remaining 8 waypoints were unreachable (displaced into the occupied space).
4 100.0 322.3 5.40 Robot stalled at Waypoints 6, 10, 18, 23, 28, 35, 37-39 due to replanning.
5 48.72 160.1 5.40 Significant map drift occurred after 2 minutes, such that the remaining 19 waypoints were unreachable as they displaced into the occupied space.
To determine the robot’s navigational performance, we evaluate it according to the two aforementioned metrics. This can be seen in Table II. In addition, in Figure  9, the time taken for each waypoint over all trials can be seen. We can see that the robot is able to reach $86.5$% of the total waypoints over all trials with a median time taken of $5.38$ s per waypoint. As seen in Figure 9, this was done in the average time of $8.525$ s with a standard deviation of $11.4$ s.
We can see several outliers where the robot took ranging from 10.0 seconds to a few minutes to complete the navigation to the adjacent waypoint, for instance, at Waypoint 17 in Trial 1. This can be explained due to Nav2 trying to create a local path that neatly ends within the desired set tolerances in position and orientation, and the fact that the robot requires a minimal input velocity in order to move, it overshoots beyond the tolerance and Nav2 has to replan the local path. This can result in the robot being stuck indefinitely, and requires assistance from a human operator. This means that for some of these outliers, ranging from 30 s, human assistance is eventually brought into action. Such that the robot is correctly placed in the desired position and orientation within the set tolerance.
Figure 9: The time taken in seconds for the robot to reach each possible unique waypoint over five trials. Gray dots describe waypoints that the robot has not been able to reach. The outliers occurred due to Nav2 getting stuck at replanning, such that the navigation terminates within narrow desired tolerances, or due to map drift.
Furthermore, we can see in Table II that in Trial 3 and 5, the robot failed to reach all of the waypoints. This is mainly due to map drifts that occurs over time and the fact that the SLAM module was still operating by mapping the robot’s surroundings online. This can be recognized e.g. as two identical hallways slightly tilted from one and another, as seen in Figure 10. Due to the drift in the navigation map, since the coordinates of the waypoints remain the same, SLAM adjusts the 2D navigation map such that the later waypoints appear in places that the robot cannot reach, in the walls for example. Drift usually occurs due to the fact that we only use relative sensors to localize itself with respect to the surroundings, e.g. IMU, leg encoders, LIDAR. This can cause the uncertainty to grow over time which is inherent in odometry.
Figure 10: A drift in the map where the hall is doubled. The drift occurred after 10 minutes after the autonomous navigation started in Trial 3.
V Conclusion
Coverage path planning allows mobile robots such as quadrupeds to explore the whole environment. This is especially useful for applications such as surveillance, inspection, and search and rescue. Nevertheless, limited work has been done on autonomous navigation using coverage path planning on quadrupeds. Therefore, we developed an open-source framework using ROS 2 that enables a Unitree Robotics Go2 quadruped to navigate such that it visits every corner using the prior 2D navigation map, autonomously. It utilizes a map reader to extract a graph of 2D waypoints using the topological skeleton of the map, and a path planner to create an efficient path with respect to time and the starting position. A state machine is used to iterate over the ordered list of waypoints and navigate them in succession.
[…]
Compared to the state-of-the-art methods that use time-constrained planning [23, 28], which do not enable variable time of task execution, our approach is not constrained based on a predefined time. While a time-constrained approach is useful when mission time is predefined, the variable-time approach provides a more adaptable solution when operating with limited knowledge. Unlike alternative state-of-the-art approaches that enable time-variable exploration [24], our approach is more computationally efficient due to the use of a morphological technique to extract the topological skeleton of the map. Even though the experimental conditions were different, a rough comparison of the map generation magnitude based on the results from each paper shows an order-of-magnitude difference (milliseconds vs microseconds).
Nevertheless, the proposed method is primarily suited for known and static environments, rendering it unsuitable for real-world applications with irregular and moving obstacles within the environment, and varying elevations. Future work will therefore focus on extending the proposed method to incorporate real-time autonomous coverage exploration in unknown 3D environments, particularly under diverse environmental conditions and in the presence of irregular and dynamic obstacles. Additionally, the influence of variations in key parameters will be systematically investigated. Moreover, the study will also investigate and seek to improve drift issues commonly encountered in SLAM. Lastly, the proposed approach will be systematically compared to existing 2D coverage path planning methods to evaluate its relative performance and advantages.